# Thought Anchors: Which LLM Reasoning Steps Matter?

### v2

@@ 1 Introduction
Training large language models to reason with chain-of-thought [28, 23, 34] has led to significant advances in capabilities [26]. The resulting reasoning traces are already used in safety research [3, 29]. However, there has been little work adapting interpretability methods to this new paradigm (though see [31, 14]). Traditional mechanistic interpretability [25, 24] methods often focus on a single forward pass of the model: understanding the activations, how they are processed by each layer, and how they are converted to the final output [33, 16, 15]. However, this framework is too fine-grained for autoregressive reasoning models, which consume their own output tokens.
A core step in many interpretability strategies is to decompose the model into smaller parts that can be analyzed independently [20]. A natural decomposition for chain-of-thought is into individual sentences and how they depend on each other. Interpretations of neural network behavior operate at varying levels of abstraction [11, 12], and sentence-level explanations strike an intermediate abstraction depth. Compared to tokens, sentences are more coherent and often coincide with reasoning steps extracted by an LLM [31, 2]. Compared to paragraphs, sentences are less likely to conflate reasoning steps and may serve as an effective target for linking different steps.
Prior work has established that different sentences within reasoning traces perform distinct functions. Backtracking sentences (e.g., “Wait…”) cause the model to revisit earlier conclusions, which boosts final-answer accuracy [22]. Other research has distinguished sentences based on whether they retrieve new information or execute deduction with existing information [31]. Hence, reasoning may follow an overarching structure, where sentences can introduce and pursue high-level computational goals. Yet, approaches for mapping this structure remain limited.
@@
First, in section 3 we present a black-box method for measuring the counterfactual impact of a sentence on the model’s final answer and future sentences. We repeatedly resample reasoning traces from the start of each sentence. Based on resampling data, we can quantify the impact of each sentence on the likelihood of any final answer or the likelihood of producing any subsequent sentence. Via this resampling approach, we can additionally distinguish planning sentences that initiate computations leading to some answer from sentences performing computations necessary for the answer but which are predetermined.
Second, in section 4 we present a white-box method for evaluating importance based on the sentences most attended. Our analyses reveal “receiver” heads that narrow attention toward particular past “broadcasting” sentences. Compared to base models, where attention is more diffuse, reasoning models display overall greater attentional narrowing through receiver heads, and these heads have an outsized impact on the model’s final answer. We develop a systematic approach to identifying receiver heads and show how evaluating sentences on the extent to which they are broadcast by these heads provides a mechanistic measure of importance.
Third and finally, in section 5 we present a method that measures the causal dependency between specific pairs of sentences in a reasoning trace. For each sentence in a trace, we intervene by masking all attention to it from subsequent tokens. We then measure the effect of suppression on subsequent token logits (KL divergence) compared to those generated during the absence of suppression. Averaging token effects by sentence, this strategy measures each sentence’s direct causal effect on each subsequent sentence.
@@ 2.1 Sentence taxonomy for reasoning functions
We categorized sentences using the framework by [31], which describes distinct reasoning functions within a reasoning trace. We define eight categories (see examples and frequencies in Appendix A):
@@
Each sentence in the analyzed response is assigned to one of these categories using an LLM-based auto-labeling approach. The prompt used to label sentence categories is provided in Appendix B. Some categories that rarely appear are omitted from the figures below. Probing classifiers achieve high accuracy in distinguishing sentences and their corresponding categories (see Appendix C).
@@ 2.2 Model and dataset
Our analyses are based on the DeepSeek R1-Distill Qwen-14B model (48 layers) [8]. We used a temperature of $0.6$ and a top-p value of $0.95$ in all of our experiments. We used the MATH dataset [17] specifically targeting 10 challenging mathematics questions that the model correctly solves 25-75% of the time, identified by testing the model on 1,000 problems 10 times each. For each selected problem, we generated one correct and one incorrect reasoning trace, producing 20 responses. The average response is 144.2 sentences long (95% CI: [116.7, 171.8] and 4208 tokens (95% CI: [3479, 4937]). We focus only on sentences before the resampling method indicates the model has converged on an answer (i.e., after which it gives the same response in >98% of resamples). In Appendix D, we provide results from applying our techniques to other reasoning models.
@@ 3.2 Counterfactual importance
    Semantic filtering. The problem with resampling importance is that in rollout resampling, if $T_{i}$ is identical or similar to $S_{i}$ then we do not get much information about whether $S_{i}$ is important or not. Therefore, we write $S\not\approx T$ if two sentences $S$ and $T$ are dissimilar, defined as having cosine similarity less than the median value across all sentence pairs in our dataset, when fed into an embeddings model (see Appendix E for details). Therefore we can finally define the counterfactual importance by conditioning on $T_{i}\not\approx S_{i}$ i.e., $\text{importance}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}}|T_{i}\not\approx S_{i})||p(A_{S_{i}})]$.
@@
Since we resample all steps after a given sentence $S_{i}$, we avoid the aforementioned limitation of forced-answering. We also provide empirical evidence that the principled counterfactual importance definition in 1-3 above is useful, by comparing it to the resampling importance in Appendix G.
@@
Figure 3: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. Only the 5 most common sentence types are shown (see Appendix F).
@@ 3.3 Sentence-to-sentence importance
Intuitively, on the right-hand side of Equation 1, the first term is the probability that a future sentence $S^{\text{Fut.}}$ will semantically occur given that $S_{i}$ was present in the trace, and the second term is the corresponding probability when $S_{i}$ is resampled with a non-equivalent sentence. A positive score indicates that sentence $S_{i}$ increases the likelihood of producing $S^{\text{Fut.}}$ (i.e., $S_{i}$ upregulates $S^{\text{Fut.}}$), while a negative score suggests that it suppresses or inhibits it.
In this context $S^{\text{Fut.}}$ semantically occurs if, when we extract the sentences and identify the best candidate match for $S^{\text{Fut.}}$ using cosine similarity between sentence embeddings, it has greater than $0.8$ cosine similarity (i.e., the median value in our dataset) to that sentence. Pseudocode for estimating sentence-to-sentence importance and empirical values of this metric can be found in Appendix G.
@@ 4.1 The identification of receiver heads
We refer to attention heads that narrow attention toward specific sentences as “receiver heads”. These heads are more common in later layers (Appendix H). To formally assess the existence of receiver heads, we tested whether some attention heads consistently operate in this role. Specifically, we measured the split-half reliability of heads’ kurtosis scores. We found a strong head-by-head correlation ($r$ = .67) between kurtosis scores computed for half of the problems with kurtosis scores for the other half of problems. Thus, some attention heads consistently operate as receiver heads, albeit with some heterogeneity across responses in which heads narrow attention most.
@@
Attentional narrowing toward particular sentences may be a feature specifically of reasoning models that enhances their performance. Comparing R1-Distill-Qwen-14B (reasoning) and Qwen-14B (base) suggests that the reasoning model’s receiver heads will narrow attention toward singular sentences to a greater degree (appendix I). Furthermore, ablating receiver heads leads to a greater reduction in accuracy than ablating self-attention heads at random (appendix J). Altogether, these findings are consistent with receiver heads and thought-anchor sentences playing special roles in reasoning.
@@ 4.2 Links to resampling importance and sentence types
[…]
@@ 5.2 Correlations with the resampling-based importance matrix
[…]
@@ 6 Case study
The three presented techniques cover different aspects of attribution within a reasoning trace. We will now illustrate the utility and the complementary nature of these techniques by applying them to the model’s response for one problem: “When the base-16 number $66666_{16}$ is written in base 2, how many base-2 digits (bits) does it have?” (Figure 6) (see Appendix L for the full transcript).
@@ 6.1 Resampling
The resampling method (section 3) reveals this initial poor trajectory and pivot upwards (shown previously in Figure 2A). Specifically, from sentences 6-12, expected accuracy steadily declines, but sentence 13 causes counterfactual accuracy to drastically increase. The pivotal role of 13 is notably missed if one instead evaluates accuracy by forcing the model to give a response immediately, as some prior studies have done, which instead yields 0
@@ 6.2 Receiver heads
The trajectory toward the final correct answer can be understood as a series of computational chunks (see flowchart in Figure 6). First, the model prepares a formula for converting $66666_{16}$ to decimal (sentences 13-19). Next, the model computes the answer to that formula, finding that $66666_{16}$ is 419,430 in decimal (sentences 20-33). The model subsequently converts that number to binary by putting forth another formula and solving it, $floor(log_{2}(419,430))+1=19$, to derive that the answer is “19 bits” (sentences 34-41). The model then notes a discrepancy with the earlier 20-bit solution (sentences 42-45). The model hence initiates new computations that verify that it computed the decimal value of $66666_{16}$ correctly (sentences 46-58) and that it computed the binary conversion accurately (sentences 59-62). Equipped with this increased certainty about 19-bit answer, the model discovers why its initial 20-bit idea was incorrect: “because leading zeros are not counted” (Sentence 66). This overall narrative is based on our analysis of attention patterns (section 4): Receiver attention heads pinpoint sentences initiating computations or stating key conclusions, thereby segmenting the reasoning trace into seemingly meaningful chunks (Figure 6).
@@ 6.3 Attention suppression
We identified these linkages based on the attention-suppression matrix (section 5), which contains local maxima at these linkages (12 → 43, 43 → 65, 12 → 66; 32 → 46, 32 → 59). Notice that many of the sentences pinpointed by the attention-suppression technique overlap with the sentences receiving high attention from receiver heads. Adding to the receiver-head conclusions, the attention suppression technique shows how information flows between these key sentences that structure the reasoning trace.
@@ 6.4 Open source interface
We released an open source interface (thought-anchors.com) for analyzing and visualizing reasoning traces. The reasoning trace can be visualized as an annotated directed acyclic graph using the importance metrics from Section 3, which may benefit interpretability and reasoning model debugging. We show our proof-of-concept interface in Figure 1C, where important sentences are represented by larger nodes and sentences that causally affect each other are connected with dashed gray lines. Hovering over a node lets users inspect sentence properties or compare alternative rollouts.
@@ 7 Related work
Chain-of-thought reasoning [34, 23, 28] has driven significant capabilities improvements in large language models. State-of-the-art performance on complex tasks comes from “thinking models” like DeepSeek-R1 [8], OpenAI’s o1 [26], and Claude 3.7 [1] are explicitly optimized for this with RLVR[36]. These increased reasoning capabilities introduce novel safety challenges, including potential unfaithfulness in the reasoning itself [19]. While current alignment evaluation methods often rely on examining reasoning traces [3, 29], we cannot rely on the model’s reasoning being faithful to the internal reasoning[7, 30, 2], creating a need for more mechanistic analyses.
Importance of individual steps. Many CoT sentences can be deleted or rewritten with limited impact on task accuracy [32, 21]. Prior attribution methods suggest that a subset of steps disproportionately drive the final answer: Shapley values [10], ROSCOE metrics [13], gradient-based scores [35], and resampling at fork tokens [4]. Complementing these, our approach provides a more principled framework for understanding how reasoning traces are constructed around key sentences.
White-box analyses of reasoning. Mechanistic studies locate specialised components that propagate information across steps: iteration or receiver heads [6], small sets of heads necessary for CoT performance [37], steerable activation directions for planning and backtracking [31], and deduction circuits in symbolic tasks [5]. Other work identifies phase shifts separating pre-training priors from contextual reasoning [9]. These findings motivate our focus on sentence-level structure rather than token-level saliency. However, they do not provide general methods for understanding the causal interdependence between sentences.
@@ 8 Discussion and Limitations
We view this as preliminary work. Our analyses require refinement to fully grapple with how downstream sentences may be overdetermined by different trajectories in a reasoning trace or independent sufficient causes. Additionally, we do not formally examine the role of error correction. Our receiver-head analyses are confounded by a sentence’s position in the reasoning trace (see Appendix M). Further, our attention-suppression work is limited because it effectively requires the model to process out-of-distribution information.

### v3

@@ 1 Introduction
Training large language models to reason with chain-of-thought (Reynolds and McDonell, 2021; Nye et al., 2021; Wei et al., 2023) has led to significant advances in capabilities (OpenAI, 2024). The resulting reasoning traces are already used in safety research (Baker et al., 2025; Shah et al., 2025). However, there has been little work adapting interpretability methods to this new paradigm (though see (Venhoff et al., 2025; Goodfire, 2025)). Traditional mechanistic interpretability (Olah et al., 2020; Olah, 2022) methods often focus on a single forward pass of the model: understanding the activations, how they are processed by each layer, and how they are converted to the final output (Wang et al., 2022; Heimersheim and Janiak, 2023; Hanna et al., 2023). However, this framework is too fine-grained for autoregressive reasoning models, which consume their own output tokens.
A core step in many interpretability strategies is to decompose the model into smaller parts that can be analyzed independently (Lindsey et al., 2025). A natural decomposition for chain-of-thought is into individual sentences and how they depend on each other. Interpretations of neural network behavior operate at varying levels of abstraction (Geiger et al., 2021, 2025), and sentence-level explanations strike an intermediate abstraction depth. Compared to tokens, sentences are more coherent and often coincide with reasoning steps extracted by an LLM (Venhoff et al., 2025; Arcuschin et al., 2025). Compared to paragraphs, sentences are less likely to conflate reasoning steps and may serve as an effective target for linking different steps.
Prior work has established that different sentences within reasoning traces perform distinct functions. Backtracking sentences (e.g., “Wait…”) cause the model to revisit earlier conclusions, which boosts final-answer accuracy (Muennighoff et al., 2025). Other research has distinguished sentences based on whether they retrieve new information or execute deduction with existing information (Venhoff et al., 2025). Hence, reasoning may follow an overarching structure, where sentences can introduce and pursue high-level computational goals. Yet, approaches for mapping this structure remain limited.
@@
First, in section˜3 we present a black-box method for measuring the counterfactual impact of a sentence on the model’s final answer and future sentences. We repeatedly resample reasoning traces from the start of each sentence. Based on resampling data, we can quantify the impact of each sentence on the likelihood of any final answer or the likelihood of producing any subsequent sentence. Via this resampling approach, we can additionally distinguish planning sentences that initiate computations leading to some answer from sentences performing computations necessary for the answer but which are predetermined.
Second, in section˜4 we present a white-box method for evaluating importance based on the sentences most attended. Our analyses reveal “receiver” heads that narrow attention toward particular past “broadcasting” sentences. Compared to base models, where attention is more diffuse, reasoning models display overall greater attentional narrowing through receiver heads, and these heads have an outsized impact on the model’s final answer. We develop a systematic approach to identifying receiver heads and show how evaluating sentences on the extent to which they are broadcast by these heads provides a mechanistic measure of importance.
Third and finally, in section˜5 we present a method that measures the causal dependency between specific pairs of sentences in a reasoning trace. For each sentence in a trace, we intervene by masking all attention to it from subsequent tokens. We then measure the effect of suppression on subsequent token logits (KL divergence) compared to those generated during the absence of suppression. Averaging token effects by sentence, this strategy measures each sentence’s direct causal effect on each subsequent sentence.
@@ 2.1 Sentence taxonomy for reasoning functions
We categorized sentences using the framework by Venhoff et al. (2025), which describes distinct reasoning functions within a reasoning trace. We define eight categories (see examples and frequencies in Appendix˜A):
@@
Each sentence in the analyzed response is assigned to one of these categories using an LLM-based auto-labeling approach. The prompt used to label sentence categories is provided in Appendix˜B. Some categories that rarely appear are omitted from the figures below. Probing classifiers achieve high accuracy in distinguishing sentences and their corresponding categories (see Appendix˜C).
@@ 2.2 Model and dataset
Our analyses are based on the DeepSeek R1-Distill Qwen-14B model (48 layers) (DeepSeek, 2025). We used a temperature of $0.6$ and a top-p value of $0.95$ in all of our experiments. We used the MATH dataset (Hendrycks et al., 2021) specifically targeting 10 challenging mathematics questions that the model correctly solves 25-75% of the time, identified by testing the model on 1,000 problems 10 times each. For each selected problem, we generated one correct and one incorrect reasoning trace, producing 20 responses. The average response is 144.2 sentences long (95% CI: [116.7, 171.8] and 4208 tokens (95% CI: [3479, 4937]). We focus only on sentences before the resampling method indicates the model has converged on an answer (i.e., after which it gives the same response in >98% of resamples). In Appendix˜D, we provide results from applying our techniques to other reasoning models.
@@ 3.2 Counterfactual importance
    Semantic filtering. The problem with resampling importance is that in rollout resampling, if $T_{i}$ is identical or similar to $S_{i}$ then we do not get much information about whether $S_{i}$ is important or not. Therefore, we write $S\not\approx T$ if two sentences $S$ and $T$ are dissimilar, defined as having cosine similarity less than the median value across all sentence pairs in our dataset, when fed into an embeddings model (see Appendix˜E for details). Therefore we can finally define the counterfactual importance by conditioning on $T_{i}\not\approx S_{i}$ i.e., $\text{importance}:=D_{\text{KL}}[p(A^{\prime}_{S_{i}}|T_{i}\not\approx S_{i})||p(A_{S_{i}})]$.
@@
Since we resample all steps after a given sentence $S_{i}$, we avoid the aforementioned limitation of forced-answering. We also provide empirical evidence that the principled counterfactual importance definition in 1-3 above is useful, by comparing it to the resampling importance in Appendix˜G.
@@
Figure 3: The mean of each sentence category for (A) forced-answer importance and (B) counterfactual importance, per the resampling method, plotted against the sentence category’s mean position in the reasoning trace. Only the 5 most common sentence types are shown (see Appendix˜F).
@@ 3.3 Sentence-to-sentence importance
Intuitively, on the right-hand side of Equation˜1, the first term is the probability that a future sentence $S^{\text{Fut.}}$ will semantically occur given that $S_{i}$ was present in the trace, and the second term is the corresponding probability when $S_{i}$ is resampled with a non-equivalent sentence. A positive score indicates that sentence $S_{i}$ increases the likelihood of producing $S^{\text{Fut.}}$ (i.e., $S_{i}$ upregulates $S^{\text{Fut.}}$), while a negative score suggests that it suppresses or inhibits it.
In this context $S^{\text{Fut.}}$ semantically occurs if, when we extract the sentences and identify the best candidate match for $S^{\text{Fut.}}$ using cosine similarity between sentence embeddings, it has greater than $0.8$ cosine similarity (i.e., the median value in our dataset) to that sentence. Pseudocode for estimating sentence-to-sentence importance and empirical values of this metric can be found in Appendix˜G.
@@ 4.1 The identification of receiver heads
We refer to attention heads that narrow attention toward specific sentences as “receiver heads”. These heads are more common in later layers (Appendix˜H). To formally assess the existence of receiver heads, we tested whether some attention heads consistently operate in this role. Specifically, we measured the split-half reliability of heads’ kurtosis scores. We found a strong head-by-head correlation ($r$ = .67) between kurtosis scores computed for half of the problems with kurtosis scores for the other half of problems. Thus, some attention heads consistently operate as receiver heads, albeit with some heterogeneity across responses in which heads narrow attention most.
@@
Attentional narrowing toward particular sentences may be a feature specifically of reasoning models that enhances their performance. Comparing R1-Distill-Qwen-14B (reasoning) and Qwen-14B (base) suggests that the reasoning model’s receiver heads will narrow attention toward singular sentences to a greater degree (appendix˜I). Furthermore, ablating receiver heads leads to a greater reduction in accuracy than ablating self-attention heads at random (appendix˜J). Altogether, these findings are consistent with receiver heads and thought-anchor sentences playing special roles in reasoning.
@@ 4.2 Links to resampling importance and sentence types
Plan generation, uncertainty management, and self checking sentences consistently receive the most attention via receiver heads (Figure 5), whereas active computation sentences receive minimal attention. Further consistent with this, sentences receiving high receiver-head attention tend to also exert a larger effect on downstream sentences per the resampling method (mean Spearman $\rho=.22$; 95% CI: [.17, .30]; detailed further in Appendix˜K). These findings are hence consistent with reasoning traces being structured around high-level sentences that initiate computations that may link high-level statements but may minimally impact the overall reasoning trajectory.
@@ 5.2 Correlations with the resampling-based importance matrix
The attention-suppression matrix values correlate with those of the resampling-method matrix. Specifically, the two matrices were positively correlated for 19 of the 20 examined reasoning traces (mean: Spearman $\rho=.20$, 95% CI: [.12, .27]). This association is stronger when considering only cases fewer than five sentences apart in the reasoning trace, which may better track direct rather than indirect effects represented by the resampling method (mean: $\rho=.34$ [.27, .40]). The magnitudes of these correlations are substantial, given that measures measure partially different aspects of causality and the resampling measure itself contains stochastic noise. Hence, these results give weight to the validity of the resampling approach, whose precision we leverage for the forthcoming case study.
@@ 6 Case study
The three presented techniques cover different aspects of attribution within a reasoning trace. We will now illustrate the utility and the complementary nature of these techniques by applying them to the model’s response for one problem: “When the base-16 number $66666_{16}$ is written in base 2, how many base-2 digits (bits) does it have?” (Figure 6) (see Appendix˜L for the full transcript).
@@ 6.1 Resampling
The resampling method (section˜3) reveals this initial poor trajectory and pivot upwards (shown previously in Figure 2A). Specifically, from sentences 6-12, expected accuracy steadily declines, but sentence 13 causes counterfactual accuracy to drastically increase. The pivotal role of 13 is notably missed if one instead evaluates accuracy by forcing the model to give a response immediately, as some prior studies have done, which instead yields 0
@@ 6.2 Receiver heads
The trajectory toward the final correct answer can be understood as a series of computational chunks (see flowchart in Figure 6). First, the model prepares a formula for converting $66666_{16}$ to decimal (sentences 13-19). Next, the model computes the answer to that formula, finding that $66666_{16}$ is 419,430 in decimal (sentences 20-33). The model subsequently converts that number to binary by putting forth another formula and solving it, $floor(log_{2}(419,430))+1=19$, to derive that the answer is “19 bits” (sentences 34-41). The model then notes a discrepancy with the earlier 20-bit solution (sentences 42-45). The model hence initiates new computations that verify that it computed the decimal value of $66666_{16}$ correctly (sentences 46-58) and that it computed the binary conversion accurately (sentences 59-62). Equipped with this increased certainty about 19-bit answer, the model discovers why its initial 20-bit idea was incorrect: “because leading zeros are not counted” (Sentence 66). This overall narrative is based on our analysis of attention patterns (section˜4): Receiver attention heads pinpoint sentences initiating computations or stating key conclusions, thereby segmenting the reasoning trace into seemingly meaningful chunks (Figure 6).
@@ 6.3 Attention suppression
We identified these linkages based on the attention-suppression matrix (section˜5), which contains local maxima at these linkages (12 → 43, 43 → 65, 12 → 66; 32 → 46, 32 → 59). Notice that many of the sentences pinpointed by the attention-suppression technique overlap with the sentences receiving high attention from receiver heads. Adding to the receiver-head conclusions, the attention suppression technique shows how information flows between these key sentences that structure the reasoning trace.
@@ 6.4 Open source interface
We released an open source interface (thought-anchors.com) for analyzing and visualizing reasoning traces. The reasoning trace can be visualized as an annotated directed acyclic graph using the importance metrics from Section˜3, which may benefit interpretability and reasoning model debugging. We show our proof-of-concept interface in Figure 1C, where important sentences are represented by larger nodes and sentences that causally affect each other are connected with dashed gray lines. Hovering over a node lets users inspect sentence properties or compare alternative rollouts.
@@ 7 Related work
Chain-of-thought reasoning (Wei et al., 2023; Nye et al., 2021; Reynolds and McDonell, 2021) has driven significant capabilities improvements in large language models. State-of-the-art performance on complex tasks comes from “thinking models” like DeepSeek-R1 (DeepSeek, 2025), OpenAI’s o1 (OpenAI, 2024), and Claude 3.7 (Anthropic, 2025) are explicitly optimized for this with RLVR(Yue et al., 2025). These increased reasoning capabilities introduce novel safety challenges, including potential unfaithfulness in the reasoning itself (Lanham et al., 2023b). While current alignment evaluation methods often rely on examining reasoning traces (Baker et al., 2025; Shah et al., 2025), we cannot rely on the model’s reasoning being faithful to the internal reasoning(Chen et al., 2025; Turpin et al., 2023; Arcuschin et al., 2025), creating a need for more mechanistic analyses.
Importance of individual steps. Many CoT sentences can be deleted or rewritten with limited impact on task accuracy (Wang et al., 2023; Madaan and Yazdanbakhsh, 2022). Prior attribution methods suggest that a subset of steps disproportionately drive the final answer: Shapley values (Gao, 2023), ROSCOE metrics (Golovneva et al., 2023), gradient-based scores (Wu et al., 2023), and resampling at fork tokens (Bigelow et al., 2024). Complementing these, our approach provides a more principled framework for understanding how reasoning traces are constructed around key sentences.
White-box analyses of reasoning. Mechanistic studies locate specialised components that propagate information across steps: iteration or receiver heads (Cabannes et al., 2024), small sets of heads necessary for CoT performance (Zheng et al., 2024), steerable activation directions for planning and backtracking (Venhoff et al., 2025), and deduction circuits in symbolic tasks (Brinkmann et al., 2024). Other work identifies phase shifts separating pre-training priors from contextual reasoning (Dutta et al., 2024). These findings motivate our focus on sentence-level structure rather than token-level saliency. However, they do not provide general methods for understanding the causal interdependence between sentences.
@@ 8 Discussion and Limitations
We view this as preliminary work. Our analyses require refinement to fully grapple with how downstream sentences may be overdetermined by different trajectories in a reasoning trace or independent sufficient causes. Additionally, we do not formally examine the role of error correction. Our receiver-head analyses are confounded by a sentence’s position in the reasoning trace (see Appendix˜M). Further, our attention-suppression work is limited because it effectively requires the model to process out-of-distribution information.